Express your answer as an ordered pair.solve by using elimination 3x-2=15x-2y=5

Answer by Edwin McCravy(20056)

    3x-2 = 15x-2y = 5

Make two equations:

         3x-2 = 15x-2y
(1)    15x-2y = 5

Simplify the first equation:

         3x-2 = 15x-2y
(2)   -12x+2y = 2

Put equations (1) and (2) together as a system to solve
by elimination:

(1)    15x-2y = 5
(2)   -12x+2y = 2

Add the two equations term by term

(1)    15x-2y = 5
(2)   -12x+2y = 2
-----------------
        3x    = 7
         x    = 7/3

Substitute 7/3 for x in (1)

(1)    15x-2y = 5
   15(7/3)-2y = 5
        35-2y = 5
          -2y = -30
            y = 15

Solution x = 7/3,  y = 15.

Edwin